OpenStudy (anonymous):

I need help with a test! Multiply or Divide. 3n^2-n2/n^2-1 divided by n^2/n+1 if possible I need the answer to be in the format similar to this one-->5(x + 2)/x + 2 * 2(n)/2(2n5) HELP! HELP HELP!

OpenStudy (swissgirl):

hmmm is it \( \large \frac{3n^2-n^2}{n^2-1} \div \frac{n^2}{n+1}\)?

OpenStudy (swissgirl):

Can you draw the function?

OpenStudy (anonymous):

Hmm I thought we could not help with tests :-(

OpenStudy (anonymous):

No that's whats the question looks like, the right answer format for it was 5(x + 2)/x + 2 * 2(n)/2(2n5) I was saying I need the answer for the top one to look like this other problems answer :)

OpenStudy (swissgirl):

Can you draw it out cuz I dont get what u r saying

OpenStudy (swissgirl):

Or Let try this

OpenStudy (swissgirl):

Is the function I typed out correct?

OpenStudy (anonymous):

To me this is what he wants. \[ \frac{\frac{3n^2-n2}{n^2-1}}{\frac{n^2}{n+1}}\]

OpenStudy (anonymous):

@MDoodler thats still looks like the question format, I need it to look this 5(x + 2)/x + 2 * 2(n)/2(2n5) only using the question with the 3n at the beginning, I need it to look typed out rather then drawn into a fraction because the question already looks that way! :/

OpenStudy (anonymous):

wait, would n+1 be the answer then? :D